OpenStudy (anonymous):

Solve the equation. (2x+1)^2=1

OpenStudy (mathstudent55):

A^2 = b where A is an expression is x and b = a non-negative number, is solved as A = sqrt(b) or A = -sqrt(b) (2x + 1)^2 = 1 becomes 2x + 1 = 1 or 2x + 1 = -1

OpenStudy (anonymous):



OpenStudy (anonymous):

wouldn't that we correct?

OpenStudy (mathstudent55):

correct